OpenStudy (anonymous):

Someone help!! find the domain and range of y=-(5-x^2)^(1/2)

OpenStudy (anonymous):

if you graph the function you can find range and domain would be all real numbers i believe im not sure though

OpenStudy (ivancsc1996):

The domain is all the real numbers, since every number will be positive due to the square, and therefore fit on the square root.

OpenStudy (ivancsc1996):

The range is for -infinity to 0 inclusive.

OpenStudy (akashdeepdeb):

We need to find the domain here and that is all the values of x that are allowed in the RHS. Now as it is under square root. x2-5 is greater than or equal to 0 because if it is lesser than 0 then the RHS becomes imaginary. So x2-5 >= 0 So domain is [sqrt 5, infiniti) Union (-infiniti, - sqrt 5] Understood? :)

OpenStudy (anonymous):

thanks guys!!! ^^

OpenStudy (anonymous):

but what about the range??

OpenStudy (anonymous):

Note that \(f(x)=-\sqrt{5-x^2}\) is the lower half of a circle with radius \(\sqrt5\) centered at the origin. This means the range is \([0,\sqrt5]\).

OpenStudy (anonymous):

Sorry, I mean \([-\sqrt5,0]\).

OpenStudy (anonymous):

Also, @AkashdeepDeb's domain doesn't seem right. For the function to be defined, you have that \(5-x^2\ge0\), or \(5\ge x^2\), which gives a solution set of \(|x|\le\sqrt5\), or \([-\sqrt5,\sqrt5]\).

OpenStudy (akashdeepdeb):

ISNT THE WHOLE THING UNDER THE SQRT?? IF NOT THEN Yeah the domain must be [-sqrt5,sqrt5] BUt is the negative sign outside the sqrt? @Mandy_Nakamoto

OpenStudy (anonymous):

yup

OpenStudy (akashdeepdeb):

Then follow her solution i took the negative sign inside the sqrt. HAHA :D

OpenStudy (anonymous):

im confused.. please explain :(

OpenStudy (anonymous):

so the domain is sqrt -5<=x=<sqrt 5 and the range is sqrt -5<=y<=0 is it? Am I right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks for the help!^^